OpenStudy (anonymous):

what is the solution set x^2+5x-6=0

OpenStudy (anonymous):

un-foil it what multiplys to get -6 and also adds to 5

OpenStudy (anonymous):

at this point if you can't find the numbers then use the quadratic equation. \[ax^2 +bx+c\] plug the variables a, b and c into the equation. \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

what are 2 numbers that, when multiplied equal to -6, but when added equal to +5? those 2 numbers are 5 and -1. so (x-1)(x+5)=0 now we have 2 solutions. x-1=0 x=1 and x+5=0 x=-5.

OpenStudy (anonymous):

not 5 and -1. Its 6 and -1.